# On importance sampling and independent Metropolis–Hastings with an unbounded weight function

### v1

@@ 5.4 An improved unbiased estimator
Consider the symmetrized UIS estimator $\tilde{F}_{u}$ in (40). Suppose that the conditions of Proposition 5.4 are satisfied. Then $\tilde{F}_{u}$ is an unbiased estimator of $\pi(f)$ for any $N\geq 1$, with finite expected cost and finite variance, and its inefficiency is equivalent to that of IS as $N\to\infty$.
The result supports the intuition that $\tilde{F}_{u}$ should be preferred to $\hat{F}_{u}$ in practice, and that the asymptotic efficiency of $\tilde{F}_{u}$ matches that of self-normalized importance sampling as $N\to\infty$, under some conditions on $\omega$ and $f$ such as those in Proposition 5.4.
@@ Acknowledgements.
Through the second author and CY Initiative Emergence, this project was supported by France 2030.
@@ A.1 Proofs of Section 1
A.1.1 Proof of (11)
We introduce the following extended state space justification of PIMH. Let the target and proposal distributions be
$\tilde{\pi}(x_{1},\ldots,x_{N},k)$ $=\frac{1}{N}\pi(x_{k})\prod_{n\neq k}q(x_{n}),$ (41)
$\tilde{q}(x_{1},\ldots,x_{N},k)$ $=\frac{\omega(x_{k})}{\sum_{n=1}^{N}\omega(x_{n})}\prod_{n=1}^{N}q(x_{n}).$ (42)
The state space is $\mathbb{X}^{N}\times\{1,\ldots,N\}$, i.e. $k$ is an integer in $\{1,\ldots,N\}$. The ratio of target divided by proposal on the extended space takes the following form:
$\tilde{\omega}(x_{1},\ldots,x_{N},k)=\frac{\tilde{\pi}(x_{1},\ldots,x_{N},k)}{\tilde{q}(x_{1},\ldots,x_{N},k)}=\frac{1}{N}\sum_{n=1}^{N}\omega(x_{n}).$ (43)
In other words, the ratio simplifies to $\hat{Z}(\mathbf{x})$, thus IMH with proposal $\tilde{q}$ and target $\tilde{q}$ is equivalent to Algorithm 3.
Draws from the proposal $\tilde{q}(x_{1},\ldots,x_{N},k)$ can be obtained by first sampling $x_{1},\ldots,x_{N}$ i.i.d. from $q$, then computing the importance weights $\omega(x_{n})$ for $n=1,\ldots,N$, and finally sampling $k$ from a Categorical distribution with probabilities given by $\omega(x_{n})/\sum_{m=1}^{M}\omega(x_{m})$. Under the extended target $\tilde{\pi}$, we have the following marginals and conditionals.
  • The marginal distribution of $k$ is the Uniform distribution on $\{1,\ldots,N\}$.
  • The marginal distribution of $x_{1},\ldots,x_{N}$ is obtained as
    $\tilde{\pi}(x_{1},\ldots,x_{N})=\sum_{n=1}^{N}\frac{1}{N}\pi(x_{n})\prod_{m\neq n}q(x_{m}).$ (44)
  • The conditional distribution of $k$ given $x_{1},\ldots,x_{N}$ can be obtained by dividing (41) by (44),
    $\tilde{\pi}(k|x_{1},\ldots,x_{N})=\frac{\omega(x_{k})}{\sum_{n=1}^{N}\omega(x_{n})}.$ (45)
  • The conditional distribution of $x_{1},\ldots,x_{N}$ given $k$ can be obtained by dividing (41) by $N^{-1}$, and is given by
    $\tilde{\pi}(x_{1},\ldots,x_{N}|k)=\pi(x_{k})\prod_{n\neq k}q(x_{n}),$ (46)
    and in particular $\tilde{\pi}(x_{k}|k)$ is $\pi(x_{k})$.
Note that $f(x_{k})$ can be written as a more explicit function of $(x_{1},\ldots,x_{N},k)$, as $\sum_{n=1}^{N}f(x_{n})\mathds{1}(n=k)$. We use the above properties to obtain the following equalities, as in Andrieu et al. (2010), Appendix B.5, either conditioning on $k$ in the first line, or on $x_{1},\ldots,x_{N}$ in the second line:
$\mathbb{E}_{\bar{\pi}}[\sum_{n=1}^{N}f(x_{n})\mathds{1}(n=k)]$ $=\mathbb{E}_{\bar{\pi}}[\mathbb{E}_{\bar{\pi}}[\sum_{n=1}^{N}f(x_{n})\mathds{1}(n=k)|k]]=\mathbb{E}_{\pi}[f],$ (47)
$=\mathbb{E}_{\bar{\pi}}[\mathbb{E}_{\bar{\pi}}[\sum_{n=1}^{N}f(x_{n})\mathds{1}(n=k)|x_{1},\ldots,x_{N}]]=\mathbb{E}_{\bar{\pi}}\left[\frac{\sum_{n=1}^{N}\omega(x_{n})f(x_{n})}{\sum_{n=1}^{N}\omega(x_{n})}\right].$ (48)
We obtain (11), i.e. the equality $\mathbb{E}_{\pi}[f]=\mathbb{E}_{\mathbf{x}\sim\bar{\pi}}[\hat{F}(\mathbf{x})]$, with $\hat{F}$ as in (4).
A.1.2 Proofs of Proposition 1.1 and Lemma 1.1
@@ A.1.1 Proof of (11)
where $m(p)$ is a positive number depending only on $p$. As described in Ren & Liang (2001), the constant $m(p)$ satisfies $(m(p))^{1/p}\leq p-1$; in fact they provide a sharper bound, but we do not need it here. For i.i.d. variables the right-hand side becomes $m(p)N^{p/2}\mathbb{E}[|X_{1}|^{p}]$. If we consider the average instead of the sum on the left, then the right-hand side becomes $m(p)N^{-p/2}\mathbb{E}[|X_{1}|^{p}]$. Since $q(\omega)=1$ and assuming that $q(\omega^{p})<\infty$, we define $X_{n}=\omega_{n}-1$ and apply the above result to obtain
@@
For the non-centred moment, we proceed as follows:
@@
This bound goes to one as $N\to\infty$. ∎
Proof of Lemma 1.1.
Using Markov’s inequality and Proposition 1.1, we have:
@@
The second line uses Markov’s inequality, and the third line applies Proposition 1.1. ∎
@@ A.2 Proofs of Section 2
Having different $x_{j}$’s in the numerator and denominator, and using their independence, the leading term in (58) is $-q(\omega^{2}\cdot(f-\pi(f)))\mathbb{E}[(\sum_{n=2}^{N}\omega(x_{n})/N)^{-2}]$.
@@
The second term leads to a bound in $N^{-s/2}$ using Markov’s inequality as in Lemma 1.1, since $q(\omega^{s})<\infty$ under the assumptions. The first term is dealt with first using Hölder’s inequality with exponents $r/s$ and $(1-s/r)^{-1}$,
@@ A.3 Proofs of Section 3
That definition only considers the probability of moves to states different than $\mathbf{x}$ that are rejected. We will use the following fact: at every iteration, for each chain one of the following three events occurs: 1) a proposal to a different state is accepted, 2) a proposal to a different state is rejected, 3) a proposal is made to the current state (and systematically accepted). In a continuous state space with an atomless measure $\lambda$, the last event occurs with probability zero.
@@ A.4 Proofs of Section 4
We then proceed through a series of lemmas to bound each term. The following lemmas are used to handle the case when $\hat{Z}(\mathbf{x})>2$:
@@
Let $\theta\in[0,1]$. We start with an $L^{p}$-version of Paley-Zygmund inequality, as in page 2705, equation (12) of Petrov (2007) with $r=1$. If $W$ is a non-negative random variable and $p>1$, then
@@
Re-arranging with $b=\theta\mathbb{E}[W]$ implies (67). We apply this to $\hat{Z}(\mathbf{x})$, under Assumption 1:
@@
For $\{\hat{Z}(\mathbf{x})\geq 1+t\}$, we directly apply Lemma 1.1:
@@
We introduce the sets $A_{k}=\left[t/(k+1),t/k\right]$ for $k\geq 1$, so that $\cup_{k=1}^{\infty}A_{k}=[0,t]$ which contains $[1,t]$. Using the result of Lemma 1.1, we obtain the bound:
@@
using Markov’s inequality as in Lemma 1.1. Finally, multiplying by $r(2)^{t-t\wedge p}$ we obtain

### v2

@@ 5.4 An improved unbiased estimator
Consider the SUIS estimator $\tilde{F}_{u}$ in (40), with cost $\tilde{\mathcal{C}}$. Suppose that the conditions of Proposition 5.4 are satisfied. Then $\tilde{F}_{u}$ is an unbiased estimator of $\pi(f)$ for any $N\geq 1$, with finite expected cost and finite variance, and its inefficiency is equivalent to that of IS as $N\to\infty$:
$\lim_{N\to\infty}\mathbb{E}[\tilde{\mathcal{C}}]\cdot\mathbb{E}\left[(\tilde{F}_{u}-\pi(f))^{2}\right]=\lim_{N\to\infty}N\cdot\mathbb{E}_{\mathbf{x}\sim\bar{q}}\left[(\hat{F}(\mathbf{x})-\pi(f))^{2}\right].$
The result supports the intuition that $\tilde{F}_{u}$ should be preferred to $\hat{F}_{u}$ in practice. Note however that the cost of the SUIS estimator is at least as large as that of the UIS estimator.
Example 6 (Examples 1-3 continued).
We perform experiments to assess the cost, variance and the inefficiency of the SUIS estimator with Exponential target and proposal distributions, and test function $f:x\mapsto\sin(x)$, as in Example 3. Figure 3 displays the variance of SUIS (left), and the product of variance times expected cost (right), as a function of $N$. For each value of $p$, as $N$ increases we observe that the variance resembles the asymptotic variance of IS divided by $2N$ represented with dashed lines (left). The inefficiency of SUIS converges to the asymptotic variance of IS represented by dashed lines (right). The results are based on $5\times 10^{6}$ independent repeats.
Figure 3: : Left: Variance of the symmetrized unbiased importance sampling estimator (solid) vs. asymptotic variance $\sigma^{2}_{\textsc{IS}}$ divided by $2N$ (dashed) across $N$, for two different values of $p$. Right: Inefficiency (variance $\times$ cost) of SUIS (solid) vs. IS (dashed).
5.5 Combining robust mean estimation with unbiased importance sampling
5.5.1 Motivation
An advantage of having eliminated the bias from IS through our SUIS approach is that we can now directly apply established robust mean estimation techniques (Lugosi & Mendelson, 2019). Their goal is to estimate the expectation $\mu$ of a random variable $X$ with finite variance $\sigma^{2}$ using $n$ i.i.d. copies $X_{1},\ldots,X_{n}$. For a given confidence level $\delta\in(0,1)$, we seek to construct an estimator $\hat{\mu}_{n}:=\hat{\mu}_{n}(X_{1},\ldots,X_{n},\delta)$ (potentially dependent on $\delta$) that satisfies, for the smallest possible value of $\epsilon=\epsilon(n,\delta)$, for all $n$:
$\mathbb{P}(|\hat{\mu}_{n}-\mu|>\epsilon)\leq\delta.$ (41)
If $\hat{\mu}_{n}$ is chosen to be the empirical average, using the Central Limit Theorem as $n\to\infty$, (41) holds asymptotically with $\epsilon=\sigma\sqrt{2\log(2/\delta)/n}$, using the inequality $\Phi^{-1}(1-\delta/2)\leq\sqrt{2\log(2/\delta)}$. On the other hand, the non-asymptotic bound obtained for the empirical mean with Chebyshev’s inequality gives $\epsilon=\sigma\sqrt{1/(n\delta)}$, which exhibits a poor dependence on $\delta$ that cannot be improved in general (see Section 2 in Lugosi & Mendelson, 2019). Remarkably, alternative, implementable estimators obtain a better dependence. The Median-of-Means (MoM, Nemirovskij & Yudin (1983)) estimator provides a prime example, where the $X_{1},\ldots,X_{n}$ are partitioned into $K$ blocks of size $m$ with $n=mK$, and the estimator is obtained as the median of the blockwise means. The following result shows that the MoM estimator achieves sub-Gaussian performance.
Theorem 5.1 (Theorem $2$ in Lugosi & Mendelson (2019)).
Let $X_{1},\ldots,X_{n}$ be i.i.d. random variables with mean $\mu\in\mathbb{R}$ and variance $\sigma^{2}\in(0,\infty)$. Let $\delta\in(0,1)$ and $K=\lceil 8\log(1/\delta)\rceil$ be a number of blocks, and assume that $n=mK$ for some positive integer $m$. Denote by $\hat{\mu}_{\textsc{MoM},n}$ the MoM estimator computed with $K$ blocks of size $m$. Then, with probability at most $\delta$, for all $n\geq\lceil 8\log(1/\delta)\rceil$,
$|\hat{\mu}_{\textsc{MoM},n}-\mu|>\sigma\sqrt{\frac{32\log(1/\delta)}{n}}.$ (42)
We provide a pseudo-code description of Median-of-Means (MoM), as well as the more efficient estimators of Minsker & Ndaoud (2021), hereafter Minsker–Ndaoud (MN), and Lee & Valiant (2022), hereafter Lee–Valiant (LV), in Appendix C.
The combination of MoM with self-normalized importance sampling (SNIS) was explored by Dau (2022). The proposed estimator, termed MoM-SNIS, is the median of $K$ SNIS estimators obtained with $M$ draws each, so that the cost is equivalent to SNIS with $N=M\times K$ draws. The above result on MoM does not directly apply because SNIS is biased for the quantity of interest. Yet Dau (2022) shows the following result.
Proposition 5.7 (Proposition 2 in Dau (2022)).
Let $\delta\in(0,1)$ and suppose that $N\geq 8(32\sigma^{2}_{\omega}\vee 1)\log(1/\delta)$, where $\sigma^{2}_{\omega}=q(\omega^{2})-1$ is the variance of the weight, assumed to be finite. Assume also that $\sigma^{2}_{\textsc{IS}}=q(\omega^{2}(f-\pi(f))^{2})$, the asymptotic variance of SNIS as in (5), is finite. Then the MoM-SNIS estimator $\hat{F}_{\textsc{MoM-SNIS}}$ with $K=\lceil 8\log(1/\delta)\rceil$ satisfies, with probability at most $\delta$,
$\left|\hat{F}_{\textsc{MoM-SNIS}}-\pi(f)\right|>\sigma_{\textsc{IS}}\sqrt{\frac{256\log(1/\delta)}{N}}$ (43)
Thus MoM-SNIS achieves sub-Gaussian performance under minimal assumptions on $\omega$ and $f$, but requires a minimum number of particles $N$ to be larger than $8(32\sigma^{2}_{\omega}\vee 1)\log(1/\delta)$, where $\sigma^{2}_{\omega}$ is typically unknown to the user. As shown in Dau (2022, Proposition 3), it is not possible to obtain a similar result for MoM-SNIS that would hold for all $N$ larger than a threshold that would depend on $\delta$ only. The combination of MN and LV with IS has not yet be studied.
In contrast, upon removing the bias of IS with SUIS, one can directly use MoM and obtain sub-Gaussian performance for all $n$ larger than $\lceil 8\log(1/\delta)\rceil$ (e.g. Theorem 5.1), where $n$ is a number of independent copies of SUIS. Indeed, using Proposition 5.1, for bounded test functions and under Assumption 2 with $p>2$, SUIS has a finite variance, for any choice of $N\geq 1$. We can similarly obtain guarantees from off-the-shelf results for MN (Minsker & Ndaoud, 2021) and LV (Lee & Valiant, 2022). Obvious disadvantages relative to MoM-SNIS (Dau, 2022) include the increased variance of SUIS relative to IS, and the random nature of the computing cost.
5.5.2 Numerical experiments in the Exponential example
We revisit the running example (Example 1-3-4), where $\pi$ is Exponential$(1)$ and $q$ is Exponential$(k)$. The weight $\omega$ has $p-\lambda$ moments for any $\lambda$ with $p=k/(k-1)$, and we consider $p\in\{2.01,2.1,3\}$. The cases represent various degrees of tail heaviness, with $p=2.01$ being the heaviest and $p=3$ the lightest. The test function $f(x)=\sin(x)$ is bounded so that SUIS has a finite variance in all cases.
We first compare the performance of self-normalized importance sampling (SNIS), exactly normalized importance sampling (ENIS) with $N=1000$, and the MoM-SNIS method of Dau (2022), tuned with $\delta=0.05$, leading to $K=\lceil 8\log(1/\delta)\rceil=24$ and we choose $M=48$ so that $K\times M=1152$ is comparable to $N$. We obtain the results shown in Table 1. The columns describe the value of $p$, the method (ENIS, SNIS or MoM-SNIS), the cost per estimator, the mean squared error (MSE), the bias, and the $95\%$ and $99\%$ quantiles of the absolute error. The results are based on $10^{6}$ independent repeats of ENIS and SNIS, and $4\times 10^{4}$ independent repeats of MoM-SNIS. Importantly, Proposition 5.7 does not apply in the cases $p\in\{2.01,2.1\}$ because $N$ is not larger than $8(32\sigma^{2}_{\omega}\vee 1)\log(1/\delta)$ in these cases. Still, the results show an advantage of MoM-SNIS for $p\in\{2.01,2.1\}$, both in terms of MSE and large quantiles of the absolute error. This comes at the cost of a noticeable bias, compared to that of SNIS. Any advantage seems to disappear in the case $p=3$.
Table 1: Performance of exactly normalized importance sampling (ENIS), self-normalized importance sampling (SNIS) and MoM-SNIS in the Exponential example.
p method cost MSE bias q95 q99
2.01 ENIS 1000 0.0023 0.0000 0.0721 0.1124
2.01 SNIS 1000 0.0021 0.0014 0.0921 0.1450
2.01 MoM-SNIS 1152 0.0016 0.0370 0.0634 0.0735
2.10 ENIS 1000 0.0018 0.0000 0.0637 0.0964
2.10 SNIS 1000 0.0016 0.0010 0.0768 0.1206
2.10 MoM-SNIS 1152 0.0015 0.0349 0.0609 0.0708
3.00 ENIS 1000 0.0004 0.0000 0.0379 0.0509
3.00 SNIS 1000 0.0004 0.0001 0.0396 0.0528
3.00 MoM-SNIS 1152 0.0006 0.0173 0.0456 0.0564
We then consider Table 2 representing the performance of MoM-SUIS, LV-SUIS, MN-SUIS, with MoM-SNIS repeated from the previous table for comparison. All methods employ $K=\lceil 8\log(1/\delta)\rceil=24$ with $\delta=0.05$, and each SUIS estimator uses $N=4$ particles. The number of estimators per block is set to $M=4$ so that the cost of each method is approximately equivalent to SNIS with $N=1000$. Theorem 5.1 does apply for MoM-SUIS for all choices of $p$, and similarly the guarantees in Lee & Valiant (2022); Minsker & Ndaoud (2021) apply to LV-SUIS and MN-SUIS. However, the numerical results suggest that MoM-SNIS performs at least as well as the other methods, despite the fact that Proposition 5.7 does not apply in the cases $p\in\{2.01,2.1\}$. All the robust mean estimation methods under consideration lead to a noticeable bias, which is expected, and to an MSE than that of SNIS in the case $p=2.01$, which is more surprising. We expect the MSE of ENIS and SNIS to eventually be the smallest as the budget $N$ increases.
Table 2: Performance of robust mean estimation techniques combined with SUIS in the Exponential example.
p method cost MSE bias q95 q99
2.01 MoM-SNIS 1152 0.0016 0.0370 0.0634 0.0735
2.01 MoM-SUIS 959 0.0018 0.0352 0.0742 0.0910
2.01 LV-SUIS 959 0.0018 0.0290 0.0717 0.0919
2.01 MN-SUIS 959 0.0017 0.0368 0.0688 0.0826
2.10 MoM-SNIS 1152 0.0015 0.0349 0.0609 0.0708
2.10 MoM-SUIS 944 0.0017 0.0343 0.0717 0.0870
2.10 LV-SUIS 944 0.0016 0.0260 0.0683 0.0873
2.10 MN-SUIS 944 0.0017 0.0372 0.0678 0.0802
3.00 MoM-SNIS 1152 0.0006 0.0173 0.0456 0.0564
3.00 MoM-SUIS 868 0.0009 0.0219 0.0556 0.0684
3.00 LV-SUIS 868 0.0006 0.0101 0.0473 0.0608
3.00 MN-SUIS 868 0.0011 0.0293 0.0562 0.0666
Finally, we visualise the quantiles of the absolute error for SNIS, MoM-SNIS, MoM-SUIS, LV-SUIS, and MN-SUIS, focusing on $p\in\{2.01,3\}$, in Figure 4. The algorithmic settings are identical to those described above. We are particularly interested in the highest quantile levels. In the case $p=2.01$, the most robust method appears to be MoM-SNIS followed by MN-SUIS, whereas in the case $p=3$, regular SNIS appears to perform best up the $99.9\%$ quantile. Thus robust mean estimation methods appear useful only in the hard cases with heavy-tailed importance weights.
Figure 4: : quantiles of the absolute error $|\hat{F}-\pi(f)|$ for SNIS, MoM-SNIS, MoM-SUIS, MN-SUIS and LV-SUIS for $p=2.01$ (top) and $p=3$ (bottom).
@@ Acknowledgements.
Through the second author and CY Initiative Emergence, this project was supported by France 2030. The fourth author acknowledges support from the National Science Foundation through grant DMS–2210849. The authors are thankful to Hai Dang Dau, Mohamed Ndaoud, Christian P. Robert and Yanbo Tang for helpful discussions.
@@ A.1 Proofs of Section 1
where $m(p)$ is a positive number depending only on $p$. As described in Ren & Liang (2001), the constant $m(p)$ satisfies $(m(p))^{1/p}\leq p-1$; in fact they provide a sharper bound, but we do not need it here. For i.i.d. variables the right-hand side becomes $m(p)N^{p/2}\mathbb{E}[|X_{1}|^{p}]$. If we consider the average instead of the sum on the left, then the right-hand side becomes $m(p)N^{-p/2}\mathbb{E}[|X_{1}|^{p}]$. Since $q(\omega)=1$ and assuming that $q(\omega^{p})<\infty$, we define $X_{n}=\omega(x_{n})-1$ and apply the above result to obtain
@@
For the non-centred moment (12), we proceed as follows:
@@
This bound gives (12), and goes to one as $N\to\infty$.
To prove (14), using Markov’s inequality and (13), we have:
@@ A.2 Proofs of Section 2
Having different $x_{j}$’s in the numerator and denominator, and using their independence, the leading term in (53) is $-q(\omega^{2}\cdot(f-\pi(f)))\mathbb{E}[(\sum_{n=2}^{N}\omega(x_{n})/N)^{-2}]$. Note that $q(|f-\pi(f)|\omega^{2})<\infty$ under the assumptions of Theorem 2.1, since $\omega^{2}<\max(\omega,\omega^{3})$ and both $q(|f-\pi(f)|\omega)<\infty$ and $q(|f-\pi(f)|\cdot\omega^{3})<\infty$.
@@
The second term leads to a bound in $N^{-s/2}$ using Markov’s inequality as in Proposition 1.1, since $q(\omega^{s})<\infty$ under the assumptions. The first term is dealt with first using Hölder’s inequality with exponents $r/s$ and $(1-s/r)^{-1}$,
@@ A.3 Proofs of Section 3
That definition only considers the probability of moves to states different than $\mathbf{x}$ that are rejected. We will use the following fact: at every iteration, for each chain one of the following three events occurs: 1) a proposal to a different state is accepted, 2) a proposal to a different state is rejected, 3) a proposal is made to the current state (and systematically accepted). In a continuous state space with an atomless measure $\lambda$, the last event occurs with probability zero. We assume Assumption 1 throughout so that $\mathbb{P}(\hat{Z}(\mathbf{x})=0)=0$ under $q$, and the states $\mathbf{x}$, $\mathbf{y}$ in this section are such that $\hat{Z}(\mathbf{x})>0$, $\hat{Z}(\mathbf{y})>0$, otherwise $r(\mathbf{x}),r(\mathbf{y})$ would not be well-defined.
@@ A.4 Proofs of Section 4
We then proceed through a series of lemmas to bound each term. The following lemmas are used to handle the case where $\hat{Z}(\mathbf{x})>2$:
@@
Let $\theta\in[0,1]$. We start with a $L^{p}$-version of the Paley-Zygmund inequality, as on page 2705, equation (12) of Petrov (2007) with $r=1$. If $W$ is a non-negative random variable and $p>1$, then
@@
Rearranging with $b=\theta\mathbb{E}[W]$ implies (62). We apply this to $\hat{Z}(\mathbf{x})$, under Assumption 1:
@@
For $\{\hat{Z}(\mathbf{x})\geq 1+t\}$, we directly apply Proposition 1.1:
@@
We introduce the sets $A_{k}=\left[t/(k+1),t/k\right]$ for $k\geq 1$, so that $\cup_{k=1}^{\infty}A_{k}=[0,t]$ which contains $[1,t]$. Using the result of Proposition 1.1, we obtain the bound:
@@
using Markov’s inequality as in Proposition 1.1. Finally, multiplying by $r(2)^{t-t\wedge p}$ we obtain